### v1

@@ 2.1 Bayesian inference without bias
Bayesian inference approaches are based on the well-known Bayes theorem
$p(\theta|\hat{\mathbf{y}})=\frac{p(\theta)p(\hat{\mathbf{y}}|\theta)}{p(\hat{\mathbf{y}})},$ (5)
@@
where $p(\theta|\hat{\mathbf{y}})$ is the posterior distribution that describes the probability of the parameter $\theta$ of having generated $\hat{\mathbf{y}}$, $p(\theta)$ represent the prior distribution based on the beliefs about $\theta$ before seeing the data, $p(\hat{\mathbf{y}}|\theta)$ is the likelihood function that represents the probability of observing $\hat{\mathbf{y}}$ for a given $\theta$ and $p(\hat{\mathbf{y}})$ is the marginal probability of observing $\hat{\mathbf{y}}$. The prior $p(\theta)$ must be prescribed, and the likelihood term $p(\hat{\mathbf{y}}|\theta)$ will consider a Gaussian error model on the residuals $\hat{\mathbf{r}}=\hat{\mathbf{y}}-f(\hat{\mathbf{x}},\theta)$ with noise $\epsilon_{\text{meas}}$. Iterative approaches such as Monte Carlo-Markov Chains (MCMC) converge to an estimation of the posterior distribution up to a constant such that $p(\theta|\hat{\mathbf{y}})\propto p(\theta)p(\hat{\mathbf{y}}|\theta)$, avoiding the calculation of $p(\hat{\mathbf{y}})$41. In this work, an MCMC algorithm with Metropolis-Hastings selection criteria19 will be applied to the parameter identification problem. This classical approach is well established and admits multiple modifications to improve its efficiency.
However, the application of classical Bayesian inference approaches presents three relevant shortcomings. First, the results obtained depend completely on the prescribed forward model $f$ and its assumptions. Modelling complex real systems requires simplifications from the geometry, governing laws or relevant parameters that cannot be quantified by an unbiased approach. The inferred parameters are intrinsic to the real system and have a meaning on their own. Due to the aforementioned simplifications, those parameters estimated through unbiased approaches will converge to values that may differ from those that would best describe the real system as a whole and present high confidence in the result. This is even more relevant in those cases where the model bias is pronounced the most. In digital twins of structures such as bridges, the simulation models tend to be extremely complex and sensors are sparse in space, which simultaneously increases the potential discrepancy between model predictions and observations due to the complexity, and the risk of overfitting due to the sparsity. Therefore, a badly tuned system that overfits to the few sensors without representing the real behaviour of the structure is an actual risk for such systems. For an infinitely growing number of uniformly distributed observation points, the posterior distribution converges to a Dirac delta distribution located at the $L_{2}$ minimizer.
Second, if we suppose an unbiased model and a prescribed measurement noise as the single source of error for the observations, any extra noise originated from other sources will not be accountable. This situation is representative of those cases where the prescribed noise is given by the sensors, as any additional noise is introduced artificially to quantify unknown effects. This introduces further errors in the inferred system. Adding a noise term as a latent parameter requires reliable prior information to inform its structure and obtain a representative posterior. Otherwise, the problem tends to be ill-posed and the noise ends up acting as a bias term with no physical meaning. If model bias is present but not accounted for in the formulation of the inference and the additional noise term is defined with zero mean, as is usually the case, this additional noise term will not be able to represent the variations due to the existing non-zero discrepancy. In structural applications, measurement noise contributes minimally to the observational variance. However, the variance produced by unobserved sources or modelling simplifications and assumptions has a big impact. This bias and variance cannot be quantified with classical unbiased approaches under prescribed noise. If the QoI is not part of the predictions but derived from them, then tuning a noise parameter would neither correct the obtained results for its value nor quantify its uncertainty.
Finally, classical unbiased models are limited in the information that they can incorporate for tuning the parameters and generating predictions. Digital twins usually have at their disposal copious amounts of data from many different sources, which can be used to enrich the model and generate more accurate predictions. However, the only way of introducing such information in these models is through a modification of the forward model, which may not always be possible. If possible, the introduction of this extra data would help to reduce the model discrepancy, indicate model deficiencies and guide potential model improvements. In a classical Bayesian setting, this presents still a challenge without a modification of the model itself.
@@ 2.2 Model bias identification approaches
The first step to correct the shortcomings of classical Bayesian methods regarding the model discrepancy error passes necessarily through acknowledging explicitly its existence. As previously mentioned, it is common practice to include a tunable noise source to absorb and quantify such model deficiencies. In cases, where the model differs greatly from the generative process, the improvements from this approach are limited. Kennedy and O’Hagan introduce in their seminal paper28 a model discrepancy formulation where the bias term $b$ is explicitly added to the computational model output for a given sample of latent parameters $\theta$:
$y(x)=f(x,\theta)+b(x)+\epsilon_{\text{meas}}(x).$ (6)
@@
In this setting, $b$ does not depend explicitly on $\theta$, as it represents the discrepancy between a given realization of the computational model and the real system, while $\epsilon_{\text{model}}$ may vary for different values of $\theta$. As the realizations of $f$ are not known a priori, $b$ must be evaluated at some point during the inference procedure. The structure of this term has to be defined beforehand as well, which will influence its behaviour during the parameter tuning. There exist two alternatives on when to evaluate $b$: after inferring the optimal parameter $\theta^{*}$ or during the inference process. Training $b$ during the parameter tuning further improves the precision and reliability of the resulting parameter $\theta^{*}$29, hence Kennedy and O’Hagan propose fitting $b$ as a Gaussian Process (GP) at every sample of $\theta$ together with the evaluation of the computational model28. The resulting final bias GP $b$ must be reconstructed from $\theta^{*}$ or an estimator of it. In this work, we will obtain the posterior distribution of $\theta$ and evaluate the Maximum Likelihood Estimator (MLE) to reconstruct $b$ from it.
We differentiate, consequently, two different responses. On the one hand, we generate the fitted response, which returns the predictions $f(x,\theta)$ for the inferred posterior distribution of $\theta$. This response is bound by the limitations of the formulation of $f$ and may not be able to represent the observations. On the other hand, we generate the bias-corrected response, which evaluates the fitted response for the MLE $\theta^{*}$, calculates the corresponding bias and returns the corrected prediction $z(x,\theta^{*})$. With a suitable GP kernel, the bias-corrected response is able to fit the observations in the absence of measurement noise. In addition, it is possible to obtain the distribution of the bias corrections for the whole posterior distribution of $\theta$.
@@
The GP for the bias $b$ is defined by a mean $\mu_{b}$ and a covariance function $C_{b}$ which depends on a kernel $k(x,x^{\prime})$. In most cases, $k(x,x^{\prime})$ represents a stationary kernel with standard deviation $\sigma_{b}$ and a length-scale $\ell_{b}$. Therefore, $b$ is defined as $b\sim GP(\mu_{b},C_{b}(\sigma_{b},\ell_{b}))$. Compared to classical Bayesian inference approaches, the only additional requirement is the need to define and train such GP for $b$. The training process of the GP would depend on the version of the framework to be implemented. This method incurs the same pitfalls of ill-posedness and extreme dependency on the prior’s choice. Nevertheless, the additional information provided by the distribution of the bias can be employed to identify deficiencies in the virtual model and its assumptions, as well as to correct the predictions of the computational model based on actual measurements.
@@
Two approaches based on Kennedy and O’Hagan’s (KOH) formulation will be implemented in the context of this work: the modularized version of KOH formulated by Bayarri et al.32 (see Section 2.2.1) and a formulation based on Orthogonal Gaussian Processes (OGPs) proposed by Plumlee40 (see Section 2.2.2). Both approaches attempt to deal with the identifiability issues that arise with Kennedy and O’Hagan’s initial formulation while preserving the predictive capabilities. In both cases, the training process is analogous (see Algorithm 1). For a given $\theta$, the training input dataset will be $\hat{\bm{x}}$ and the training output will be the corresponding residuals evaluated as the difference between the measurements and the computational model response $\hat{\bm{r}}=\hat{\bm{y}}-f(\theta,\hat{\bm{x}})$. The parameters of $b$ will be obtained from its maximum a posteriori probability (MAP) estimate. The difference in both methods resides in the definition of the kernels and the covariance matrix of the GP. To our knowledge, they are yet to be applied and evaluated in the context of large structures or simulation-based digital twins. Additionally, these implementations are further extended to identify the noise in biased models and to enhance the model through an extension of the bias’ input domain, tackling two of the main shortcomings of classical Bayesian approaches.
@@
2:     for Evaluation of $\theta$ in the MCMC run sampled from the last state of the chain do
3:         Define the bias model $b(x)$ as a GP on the residuals $\hat{\mathbf{r}}=\hat{\mathbf{y}}-f(\hat{\mathbf{x}},\theta)$ for a given $\theta$
@@
6:         Use the marginal likelihood of $b$ and the prior of $\theta$ to obtain the posterior distribution for the sampled $\theta$ as in Equation 8
7:         Decide if keeping or rejecting $\theta$ based on its posterior
@@
10:Regenerate the full bias-corrected response $f(x,\theta^{*})+b_{\theta}(x)$ for the estimated optimum value $\theta^{*}$.
2.2.1 Modularized Kennedy and O’Hagan’s formulation
In the original KOH’s formulation, the GPs for the bias and computational models are trained simultaneously with the inference of $\theta$. If any of the components are badly defined due to deficient priors or other error sources, the whole process gets degraded32. The contamination between different components can be avoided by modularizing the algorithm. This technique is common practice in engineering systems, and results in a better-defined and less intrusive methodology. A key distinction is the use of the marginal log-likelihood of the fitted GP of $b$ for a given $\theta$ as the marginal log-likelihood of the sampled $\theta$ itself. As $b$ is defined as a GP, its marginal log-likelihood can be obtained as:
$\mathcal{L}(\hat{\mathbf{x}},\theta)=-\frac{1}{2}\hat{\mathbf{r}}^{\top}(C_{b}+\sigma_{n}^{2}\mathbf{I})^{-1}\hat{\mathbf{r}}-\frac{1}{2}\log|{C_{b}+\sigma_{n}^{2}\mathbf{I}}|-\frac{n}{2}\log 2\pi$ (7)
where the residuals $\hat{\mathbf{r}}$ are defined as $\hat{\mathbf{r}}=\hat{\mathbf{y}}-f(\hat{\mathbf{x}},\theta)$, and $\sigma_{n}^{2}\mathbf{I}$ is the covariance generated by the prescribed measurement noise with amplitude $\sigma_{n}$ associated with $\epsilon_{\text{meas}}$. This value is computed using the algorithm 2.1 from Rasmussen42. Therefore, the posterior distribution fulfils
$p(\theta|\hat{\mathbf{y}})\propto p(\theta)p(b(\hat{\mathbf{x}})|\theta)=p(\theta)\exp(\mathcal{L}(\hat{\mathbf{x}},\theta)).$ (8)
This approach is the current standard to apply Kennedy and O’Hagan’s methodology due to its simple implementation43. For those approaches that include a bias term, two different sets of results can be observed. On the one hand, we can generate the posterior predictive distribution for the system as it was done for the case with no bias. As previously mentioned, considering the likelihood of the GP for the bias when evaluating a candidate for $\theta$ adds information on the whole generative process, which prevents the posterior predictive from converging to the MSE with respect to the observations. This awareness of the underlying generative process makes the prediction less sensitive to model discrepancy.
@@ 2.2.2 Orthogonal Gaussian Processes
To mitigate the identifiability problem of approaches based on KOH’s formulation, Plumlee proposes modelling the discrepancy term as an Orthogonal Gaussian Process (OGP) 40. This is achieved by imposing an orthogonality restriction over the bias such that it represents exclusively the discrepancy between observations and predictions that could not be explained by $f$, supposing that the given sampled values for $\theta$ were the best-fitting ones. Therefore, the bias is defined as orthogonal to the gradient of a specific loss function, $L^{2}$ loss in this case, with respect to the latent parameters.
@@
$C_{b_{\theta}}(x,x^{\prime})=k_{b}(x,x^{\prime})-h_{\theta}(x)^{T}H_{\theta}^{-1}h_{\theta}(x^{\prime})$ (9)
@@
$h_{\theta}(x)=\dfrac{1}{N}\sum_{i=1}^{N}\dfrac{\partial}{\partial\theta}f(\bm{\xi}_{i},\theta)k_{b}(x,\bm{\xi}_{i})$ (10)
@@
$H_{\theta}=\dfrac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\dfrac{\partial}{\partial\theta}f(\bm{\xi}_{i},\theta)\left[\dfrac{\partial}{\partial\theta}f(\bm{\xi}_{i},\theta)\right]^{T}k_{b}(\bm{\xi}_{j},\bm{\xi}_{i}),$ (11)
where $k_{b}$ is the original correlation function prescribed by a kernel and a standard deviation. The anchor points can be defined, for example, as the nodes of a FEM simulation, the training points of a GP or just a set of points where $f$ must be evaluated as a discrete approximation of $f$ to determine the orthogonality condition. In contrast with the other implemented approach, these anchor points do not necessarily coincide with the observation points. Therefore, the definition of $\hat{\bm{\xi}}$ introduces a potential difference between the MSE of the observations and the inferred optimum for the $L^{2}$-norm. These equations can be expressed in matrix form, being $t$ the dimensionality of $\theta$, as
$C_{b_{\theta}}(x,x^{\prime})=k_{b}(x,x^{\prime})-w(x)^{T}F_{\theta}\left(F_{\theta}^{T}WF_{\theta}\right)^{-1}F_{\theta}^{T}w(x^{\prime})$ (12)
where $w(x)$ is the $N\times 1$ vector with elements $k_{b}(x,\bm{\xi}_{i})$, $F_{\theta}$ is the $N\times t$ with rows $\dfrac{\partial}{\partial\theta}f(\bm{\xi}_{i},\theta)$ and $W$ is an $N\times N$ matrix with elements $k_{b}(\bm{\xi}_{j},\bm{\xi}_{i})$.
The introduction of the orthogonality condition mitigates the identifiability issue without requiring a well-defined prior distribution and reduces the variability of the inferred latent parameters40. These are valuable advantages in the case of digital twins, where the increased computational costs is only relevant for the off-line phase of the model, where the focus lies on the calibration and update of the model and not on its fast evaluation through new data points.
The inferred parameters with OGP bias tend to converge to the optimal value for the predefined loss and present narrower distributions than the modularized formulation of KOH due to the orthogonality condition. This effect is especially relevant if the optimum value of the latent parameters over the whole system domain is expected to differ largely from the optimum over only the sampled points. However, in contrast to KOH’s implementation and its modularized version, OGPs require the evaluation of the derivatives of $f$ with respect to the latent parameters $\theta$, which increases its evaluation times.

### v2

@@ 2.1 Bayesian inference without bias
Bayesian inference approaches are based on Bayes theorem44
$p(\bm{\theta}|\hat{\mathbf{y}})=\frac{p(\bm{\theta})p(\hat{\mathbf{y}}|\bm{\theta})}{p(\hat{\mathbf{y}})},$ (6)
@@
where $p(\bm{\theta}|\hat{\mathbf{y}})$ is the posterior distribution that describes the probability of the parameters $\bm{\theta}$ of having generated $\hat{\mathbf{y}}$, $p(\bm{\theta})$ represent the prior distribution based on the beliefs about $\theta$ before seeing the data, $p(\hat{\mathbf{y}}|\theta)$ is the likelihood function that represents the probability of observing $\hat{\mathbf{y}}$ for a given $\theta$ and $p(\hat{\mathbf{y}})$ is the marginal probability of observing $\hat{\mathbf{y}}$. The prior $p(\bm{\theta})$ must be prescribed, and the likelihood term $p(\hat{\mathbf{y}}|\theta)$ will consider a Gaussian error model on the residuals $\hat{\mathbf{r}}=\hat{\mathbf{y}}-f(\hat{\mathbf{x}},\theta)$ with noise $\epsilon_{\text{meas}}$. Iterative approaches such as Monte Carlo-Markov Chains (MCMC) converge to an estimation of the posterior distribution up to a constant such that $p(\bm{\theta}|\hat{\mathbf{y}})\propto p(\bm{\theta})p(\hat{\mathbf{y}}|\theta)$, avoiding the calculation of $p(\hat{\mathbf{y}})$45. In this work, an MCMC algorithm with Metropolis-Hastings selection criteria44 will be applied to the parameter identification problem. This classical approach is well established and admits multiple modifications to improve its efficiency. We differentiate between unbiased models, where $\epsilon_{\text{model}}=0$ is assumed for Equation 2 and biased models, where an additive discrepancy term $\epsilon_{\text{model}}\neq 0$ is included. Then, it is necessary to specify a model to quantify $\epsilon_{\text{model}}$ and its possible dependency on $\mathbf{x}$.
Within this paper, the parameters $\bm{\lambda}$ will not be explicitly included in the formulation to avoid excessive notation, as they do not play a role in the analysis that will be presented. These are parameters that, despite generally being relevant for the computational model’s formulation and evaluation, are prescribed beforehand and are neither inferred nor used as variable input.
With the generalization of monitoring systems and the ageing of bridges, the necessity to tune a simulation model that represents the bridge behaviour has increased. In most common applications, the available observations correspond to acceleration measurements of a set of sensors dispose over the bridge, coupled with condition sensors such as temperature and humidity13. Then, a dynamic structural model of the bridge is fitted such that the harmonic response obtained from the acceleration corresponds with the model’s one 10, 43. Therefore, the vector of parameters $\bm{\theta}$ corresponds with the geometrical and material properties of the bridge, e.g. Young’s modulus $E$, and the target outputs are the eigenfrequencies and associated eigenvectors. Although in some recent cases the model discrepancy has been included in the parameter inference46, 47, the common practice is to increase the complexity of the model until there is a good agreement between predictions and observations10, 43, 48 either through frequentist or Bayesian inference approaches.
@@ 2.2 Model bias identification approaches
The first step to correct the shortcomings of classical Bayesian methods regarding the model discrepancy error passes necessarily through acknowledging explicitly its existence. As previously mentioned, it is common practice to include a tunable noise source to absorb and quantify such model deficiencies. In cases, where the model differs greatly from the generative process, the improvements from this approach are limited. Kennedy and O’Hagan introduce in their seminal paper27 a model discrepancy formulation where the bias term $b$ is explicitly added to the computational model output for a given sample of latent parameters $\bm{\theta}$:
$y(x)=f(x,\bm{\theta})+b(x)+\epsilon_{\text{meas}}(x).$ (7)
@@
As the realizations of $f$ are not known a priori, $b$ must be evaluated at some point during the inference procedure. The structure of this term has to be defined beforehand as well, which will influence its behaviour during the parameter tuning. There exist two alternatives on when to evaluate $b$: after inferring the optimal parameter $\theta^{*}$ or during the inference process. Training $b$ during the parameter tuning further improves the precision and reliability of the resulting parameter $\bm{\theta}^{*}$28, hence Kennedy and O’Hagan propose fitting $b$ as a Gaussian Process (GP) at every sample of $\theta$ together with the evaluation of the computational model27.
The GP for the bias $b$ is defined by a mean $\mu_{b}$ and a covariance function $C_{b}$ which depends on a kernel $k(x,x^{\prime})$. In most cases, $k(x,x^{\prime})$ represents a stationary kernel with standard deviation $\sigma_{b}$ and a length-scale $\ell_{b}$. Therefore, $b$ is defined as
$b\sim GP(\mu_{b},C_{b}(\sigma_{b},\ell_{b})).$ (8)
We differentiate, consequently, two different responses. On the one hand, we generate the fitted response, which returns the predictions $f(x,\bm{\theta})+\epsilon_{\text{meas}}$ for the inferred posterior distribution of $\bm{\theta}$. This fitted response is a random variable that we will characterize by its mean and variance as
$\mu^{f}=\mathbb{E}\left[f(x,\bm{\theta})+\epsilon_{\text{meas}}|\hat{\mathbf{y}}\right]$ (9)
$(\sigma^{f})^{2}=\text{Var}\left(f(x,\bm{\theta})+\epsilon_{\text{meas}}|\hat{\mathbf{y}}\right).$ (10)
This response is bounded by the limitations of the formulation of $f$ and may not be able to represent the observations. On the other hand, we generate the bias-corrected response as the output where the bias has been added as,
$\mu_{\text{dist}}^{h}=\mathbb{E}\left[f(x,\bm{\theta})+b_{\theta}(x)+\epsilon_{\text{meas}}|\hat{\mathbf{y}}\right]$ (11)
$(\sigma_{\text{dist}}^{h})^{2}=\text{Var}\left(f(x,\bm{\theta})+b_{\bm{\theta}}(x)+\epsilon_{\text{meas}}|\hat{\mathbf{y}}\right),$ (12)
where the bias term has been explicitly indicated as $b_{\bm{\theta}}(x)$ to highlight that each realization depends on the value of $\bm{\theta}$. This correlation of the GP of $b$ and $\bm{\theta}$ presents a caveat for the fully Bayesian setting, as calculating $p(\hat{\mathbf{y}}|\bm{\theta})$ requires marginalizing the GP. Additionally, it is often not practical to evaluate $\mu^{h}$ and $\sigma^{h}$ depending on the full distribution of $\bm{\theta}$. Therefore, for practical reasons, the calculation of $b$ will be obtained using an estimation step. Taking inspiration in previous works31, we will use a Maximum A Posteriori (MAP) estimation of $\bm{\theta}$ as
$\bm{\theta}^{*}=\underset{\bm{\theta}}{\arg}\max p(\bm{\theta})p(\hat{\mathbf{y}}|\bm{\theta})$ (13)
to calculate the responses that include bias:
$\mu^{h}=\mathbb{E}\left[f(x,\bm{\theta}^{*})+b_{\bm{\theta}^{*}}(x)+\epsilon_{\text{meas}}|\hat{\mathbf{y}}\right]$ (14)
$(\sigma^{h})^{2}=\text{Var}\left(f(x,\bm{\theta}^{*})+b_{\bm{\theta}^{*}}(x)+\epsilon_{\text{meas}}|\hat{\mathbf{y}}\right).$ (15)
The resulting final bias GP $b_{\bm{\theta}^{*}}(x)$ must then be reconstructed from $\bm{\theta}^{*}$.
@@
Compared to classical Bayesian inference approaches, the only additional requirement is the need to define and train such GP for $b$. The training process of the GP would depend on the version of the framework to be implemented. This method incurs the same pitfalls of ill-posedness and extreme dependency on the prior’s choice. Nevertheless, the additional information provided by the distribution of the bias can be employed to identify deficiencies in the virtual model and its assumptions, as well as to correct the predictions of the computational model based on actual measurements.
@@
Two approaches based on Kennedy and O’Hagan’s (KOH) formulation will be implemented in the context of this work: the modularized version of KOH formulated by Bayarri et al.31 (see Section 2.2.1) and a formulation based on Orthogonal Gaussian Processes (OGPs) proposed by Plumlee42 (see Section 2.2.2). Both approaches attempt to deal with the identifiability issues that arise with Kennedy and O’Hagan’s initial formulation while preserving the predictive capabilities. In both cases, the training process is analogous (see Algorithm 1). For a given $\bm{\theta}$, the training input dataset will be $\hat{\bm{x}}$ and the training output will be the corresponding residuals evaluated as the difference between the measurements and the computational model response $\hat{\bm{r}}=\hat{\bm{y}}-f(\bm{\theta},\hat{\bm{x}})$. The parameters of $b$ will be obtained from its maximum a posteriori probability (MAP) estimate as
$\mu_{b}^{*},\sigma_{b}^{*},\ell_{b}^{*}=\underset{\bm{\theta}}{\arg}\max p(\hat{\bm{r}}|\mu_{b},\sigma_{b},\ell_{b})p(\mu_{b},\sigma_{b},\ell_{b}),$ (16)
where $p(\hat{\bm{r}}|\mu_{b},\sigma_{b},\ell_{b})$ returns the probability of $b(x,\mu_{b},\sigma_{b},\ell_{b})$ of representing $\hat{\bm{r}}$. This is a well established approach of fitting $b$49, although a fully Bayesian approach with Gibbs sampling alternating between samples of $\bm{\theta}$ and $(\mu_{b},\sigma_{b},\ell_{b})$ would be possible but computationally more demanding.
The difference in both methods resides in the definition of the kernels and the covariance matrix of the GP. To our knowledge, they are yet to be applied and evaluated in the context of large structures or simulation-based digital twins. Additionally, these implementations are further extended to identify the noise in biased models and to enhance the model through an extension of the bias’ input domain, tackling two of the main shortcomings of classical Bayesian approaches.
2.2.1 Modularized Kennedy and O’Hagan’s formulation
In the original KOH’s formulation, the GPs for the bias and computational models are trained simultaneously with the inference of $\bm{\theta}$. If any of the components are badly defined due to deficient priors or other error sources, the whole process gets degraded31. The contamination between different components can be avoided by modularizing the algorithm. This technique is common practice in engineering systems, and results in a better-defined and less intrusive methodology. A key distinction is the use of the marginal log-likelihood of the fitted GP of $b$ for a given $\bm{\theta}$ as the marginal log-likelihood of the sampled $\bm{\theta}$ itself. As $b$ is defined as a GP, its marginal log-likelihood can be obtained as:
$\mathcal{L}(\hat{\mathbf{x}},\bm{\theta})=-\frac{1}{2}\hat{\mathbf{r}}^{\top}(C_{b}+\sigma_{n}^{2}\mathbf{I})^{-1}\hat{\mathbf{r}}-\frac{1}{2}\log|{C_{b}+\sigma_{n}^{2}\mathbf{I}}|-\frac{n}{2}\log 2\pi$ (17)
where the residuals $\hat{\mathbf{r}}$ are defined as $\hat{\mathbf{r}}=\hat{\mathbf{y}}-f(\hat{\mathbf{x}},\bm{\theta})$, and $\sigma_{n}^{2}\mathbf{I}$ is the covariance generated by the prescribed measurement noise with amplitude $\sigma_{n}$ associated with $\epsilon_{\text{meas}}$. This value is computed using the algorithm 2.1 from Rasmussen50. Therefore, the posterior distribution fulfils
$p(\bm{\theta}|\hat{\mathbf{y}})\propto p(b(\hat{\mathbf{x}})|\bm{\theta})p(\bm{\theta})=\exp(\mathcal{L}(\hat{\mathbf{x}},\bm{\theta}))p(\bm{\theta}).$ (18)
This approach is the current standard to apply Kennedy and O’Hagan’s methodology due to its simple implementation49. For those approaches that include a bias term, two different sets of results can be observed. We can generate the posterior predictive distribution for the system as it was done for the case with no bias. As previously mentioned, considering the likelihood of the GP for the bias when evaluating a candidate for $\bm{\theta}$ adds information on the whole generative process, which prevents the posterior predictive from converging to the MSE with respect to the observations. This awareness of the underlying generative process makes the prediction less sensitive to model discrepancy.
@@ 2.2.1 Modularized Kennedy and O’Hagan’s formulation
2:     for Evaluation of $\bm{\theta}$ in the MCMC run sampled from the last state of the chain do
3:         Define the bias model $b(x)$ as a GP on the residuals $\hat{\mathbf{r}}=\hat{\mathbf{y}}-f(\hat{\mathbf{x}},\bm{\theta})$ for a given $\bm{\theta}$
@@
6:         Use the marginal likelihood of $b$ and the prior of $\bm{\theta}$ to obtain the posterior distribution for the sampled $\bm{\theta}$ as in Equation 18
7:         Decide if keeping or rejecting $\bm{\theta}$ based on its prior and likelihood evaluations
@@
10:Regenerate the full bias-corrected response $f(x,\bm{\theta}^{*})+b(x)$ for the estimated optimum value $\bm{\theta}^{*}$.
@@ 2.2.2 Orthogonal Gaussian Processes
To mitigate the identifiability problem of approaches based on KOH’s formulation, Plumlee proposes modelling the discrepancy term as an Orthogonal Gaussian Process (OGP) 42. This is achieved by imposing an orthogonality restriction over the bias such that it represents exclusively the discrepancy between observations and predictions that could not be explained by $f$, supposing that the given sampled values for $\bm{\theta}$ were the best-fitting ones. Therefore, the bias is defined as orthogonal to the gradient of a specific loss function, $L^{2}$ loss in this case, with respect to the latent parameters.
@@
$C_{b_{\theta}}(x,x^{\prime})=k_{b}(x,x^{\prime})-\bm{h}_{\theta}(x)^{T}H_{\theta}^{-1}\bm{h}_{\theta}(x^{\prime})$ (19)
@@
$\bm{h}_{\theta}(x)=\dfrac{1}{N}\sum_{i=1}^{N}\dfrac{\partial}{\partial\bm{\theta}}f(\bm{\xi}_{i},\bm{\theta})k_{b}(x,\bm{\xi}_{i})$ (20)
@@
$H_{\theta}=\dfrac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\dfrac{\partial}{\partial\bm{\theta}}f(\bm{\xi}_{i},\theta)\left[\dfrac{\partial}{\partial\bm{\theta}}f(\bm{\xi}_{i},\bm{\theta})\right]^{T}k_{b}(\bm{\xi}_{j},\bm{\xi}_{i}),$ (21)
where $k_{b}$ is the original correlation function prescribed by a kernel and a standard deviation. A subindex $\theta$ has been added, following the original implementation from Plumlee42, to emphasise that $C_{b_{\theta}}$, $\bm{h}_{\theta}(x)$ and $H_{\theta}$ are only valid for the sampled $\bm{\theta}$ at which they are evaluated. The anchor points can be defined, for example, as the nodes of a mesh that discretizes the geometry of the system, the training points of a GP or just a set of points where $f$ must be evaluated as a discrete approximation of $f$ to determine the orthogonality condition. In summary, the anchor points correspond to those input values that define the system’s response in the domain of interest. In contrast with the other implemented approach, these anchor points do not necessarily coincide with the observation points. For example, given a simulation of a physical system whose geometry is modelled with a discrete mesh, the observation points will be those for which there are measurements, while the anchor points will be all the nodes of the mesh. Therefore, the definition of $\hat{\bm{\xi}}$ introduces a potential difference between the MSE of the observations and the inferred optimum for the $L^{2}$-norm, which enforces the orthogonality condition over $\hat{\bm{\xi}}$ and not only over the observations. These equations can be expressed in matrix form, being $t$ the dimensionality of $\bm{\theta}$, as
$C_{b_{\theta}}(x,x^{\prime})=k_{b}(x,x^{\prime})-\bm{w}(x)^{T}F_{\theta}\left(F_{\theta}^{T}WF_{\theta}\right)^{-1}F_{\theta}^{T}\bm{w}(x^{\prime})$ (22)
where $\bm{w}(x)$ is the $N\times 1$ vector with elements $k_{b}(x,\bm{\xi}_{i})$, $F_{\theta}$ is the $N\times t$ matrix with rows $\dfrac{\partial}{\partial\bm{\theta}}f(\bm{\xi}_{i},\bm{\theta})$ and $W$ is an $N\times N$ matrix with elements $k_{b}(\bm{\xi}_{j},\bm{\xi}_{i})$.
The introduction of the orthogonality condition mitigates the identifiability issue without requiring a well-defined prior distribution and reduces the variability of the inferred latent parameters42. These are valuable advantages in the case of digital twins, where the increased computational costs is only relevant for the off-line phase of the model, where the focus lies on the calibration and update of the model and not on its fast evaluation through new data points.
The inferred parameters with OGP bias tend to converge to the optimal value for the predefined loss and present narrower distributions than the modularized formulation of KOH due to the orthogonality condition. This effect is especially relevant if the optimum value of the latent parameters over the whole system domain is expected to differ largely from the optimum over only the sampled points. However, in contrast to KOH’s implementation and its modularized version, OGPs require the evaluation of the derivatives of $f$ with respect to the latent parameters $\bm{\theta}$, which increases its evaluation times.